OpenStudy (camerondoherty):

LaToya decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a beginner's fee of $100 and a monthly fee of $38. If x represents the number of months that LaToya is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time: 100 + 38x = C LaToya has allocated a maximum of $404 to spend on her gym membership. Which number line shows the possible number of months that LaToya can be a member of the gym?

OpenStudy (anonymous):

maybe 5

OpenStudy (camerondoherty):

Can i brb for lunch?

OpenStudy (here_to_help15):

LaToya decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a beginner's fee of $100 and a monthly fee of $38. If x represents the number of months that LaToya is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time:

OpenStudy (here_to_help15):

So u know @camerondoherty i copied and pasted that website is https://answers.yahoo.com/question/index?qid=20110812130358AA9Bbnr

OpenStudy (camerondoherty):

Oh ok... but which one is it?

OpenStudy (camerondoherty):

do u have an equation?

OpenStudy (here_to_help15):

this is an easy one. 100+38x=404 100-100+38x=404-100 38x=303 38/38x=304/38 x=8

OpenStudy (here_to_help15):

thats what i meant to paste lol

OpenStudy (camerondoherty):

i dont think its supposed to be = sign...

OpenStudy (here_to_help15):

well hmmm now im stumped

OpenStudy (camerondoherty):

bbl

OpenStudy (here_to_help15):

ok later cam

OpenStudy (here_to_help15):

i meant cameron thats what i call my friend for short :p

OpenStudy (here_to_help15):

@camerondoherty